OpenStudy (anonymous):

does anyone know how to do geometric sequences with explicit formulas & recrusive formulas? i dont get it at all ._.

OpenStudy (anonymous):

a recursive formula would be: \[a_{n+1} = a_nr\] where r is the common ratio between the terms. For example: 1, 2, 4, 8, 16, 32,.... In this sequence, the common ratio is 2, and to get the next term in the sequence, we multiply the current term by 2.

OpenStudy (anonymous):

For a closed formula, you can figure it out as such. Lets say i have an initial term and a common ratio: \[a_0, r\] then the next term would be: \[a_1 = a_0r\] and the next term would be: \[a_2 = a_1r = a_0r^2\] and the next: \[a_3 = a_2r = a_1r^2 = a_0r^3\] and so on, so a closed formula would be: \[a_n = a_0r^n\] Now be a little careful. Since we started the sequence at the "zero-th" term, if the question asks for the 6th term, that means you need: \[a_5 = a_0r^5\]

OpenStudy (anonymous):

but how do you know which numbers to plug into where?

OpenStudy (anonymous):

you always what to know what your first term and your common ratio is. from there you can figure anything out. For example: 7, 14, 28, 56, 112,.... First i ask, "What is the first term?" Thats kinda easy, its 7. Then I ask, "what is the common ratio?" To figure out, just divide two terms next to each other. You could do 14/7 = 2, or 112/56 = 2. It doesnt matter. So now you have your common ratio. Now you can answer anything. What is the recursive formula for this sequence? \[a_{n+1} = 2a_n\] What is the closed forumla for this sequence? \[a_n = a_0r^n \Rightarrow a_n = 7(2)^n\] What is the 7th term? \[a_6 = 7(2)^6 \Rightarrow a_6 = 7*64 \Rightarrow a_6 = 448\] Remember, because i start the sequence at 0, the 7th term is a6, NOT a7